Grade 10General Physics

A body falls freely from the top of a tower . It covers 36% of the total height in the last second before striking the ground level. Find the height of tower

Moustafa Youssef Abdallah Nassar
7 Years agoGrade 10

1 Answer

Arun
7 Years ago
Dear student
 
Let H be the height of the building and t be the time to fall through this height H. 

Then H = ½ g t^2 = 5 t^2. ----------------- (1) 

In (t-1) second the distance traveled is 1 - 0.36 H = 0.64 H. 

0.64 H = 5 (t-1)^2.---------------------------(2) 

(2) / (1) gives 0.64 = (t-1)^2/ t^2 

Or 0.8 = (t-1) / t 

Solving we get t= 5 second. 

Using (1) we get H = 125 m.
 
hope it helps